OpenStudy (andrewkaiser333):

A cup of gold colored metal beads was measured to have a mass of 425g. By water displacement, the volume of beads was calculated to be 48.0cm^3. Given the following densities, identify the metal. Gold: 19.3 g/ml Copper: 8.86 g/ml Bronze: 9.87 g/ml

OpenStudy (andrewkaiser333):

@MikeZack123

OpenStudy (andrewkaiser333):

@Dark_Riddles @dan815

OpenStudy (andrewkaiser333):

@Ashleyisakitty

OpenStudy (koikkara):

Alright, Its like you have tagged all professionals, Please don't spam Chat, Wait in patience..!

OpenStudy (andrewkaiser333):

well if no one answers what am i supposed to do? i have no teacher helping i am stressed out and having troubles with health so i am trying to figure out this thing

OpenStudy (koikkara):

You can say it in the chat no one is helping, and soon someone will surely get to you, hope that is much simple..

OpenStudy (andrewkaiser333):

am sorry for it but i cannot do anything else till i understand this one to do the others the others are exactally the same

OpenStudy (koikkara):

alright, so from question, what do you want to find out ? And what are the given data ?

OpenStudy (koikkara):

Hmm 1 cm3 of water equals 1 mL. Hope YOU know it ! Then what can You Do ?

OpenStudy (koikkara):

Well, calculating I found copper looks quite good ! and what do you get ?

OpenStudy (koikkara):

@AndrewKaiser333 Can You follow me ?

OpenStudy (andrewkaiser333):

sorry was using rest room

OpenStudy (andrewkaiser333):

that is the thing it is confusing me can you go step by step.

OpenStudy (koikkara):

Alright again don't goo to Bedrooms, Stay here... So whats the formula to calculate Density ?

OpenStudy (andrewkaiser333):

d=m/v

OpenStudy (koikkara):

Good, So what Mass is given in question ?

OpenStudy (andrewkaiser333):

m= mass v= volume d= density

OpenStudy (andrewkaiser333):

425g

OpenStudy (koikkara):

Yeah, you are right.... volume ?

OpenStudy (andrewkaiser333):

48.0cm^3

OpenStudy (andrewkaiser333):

?

OpenStudy (koikkara):

Wow amazing, so as you said D= M/V which is D= 425g/48.0cm3 = ?

OpenStudy (andrewkaiser333):

8.85g/cm^3?

OpenStudy (koikkara):

Sure ? So which option is that ?

OpenStudy (andrewkaiser333):

but then it says "Given the following densities, identify the metal." is that for Gold: 19.3 g/ml Copper: 8.86 g/ml Bronze: 9.87 g/ml ????

OpenStudy (koikkara):

While we take values, look at the closer values to your solution....i mean the nearest one....... And so You did it all by You, !! Incredible Work ! Nice to meet You ! Have A Nice Day !

OpenStudy (koikkara):

Finally which one is close to your solution ?

OpenStudy (koikkara):

@AndrewKaiser333 Hello, again wash room ?

OpenStudy (andrewkaiser333):

but then it says "Given the following densities, identify the metal." is that for Gold: 19.3 g/ml Copper: 8.86 g/ml Bronze: 9.87 g/ml ????

OpenStudy (andrewkaiser333):

what do i do with those????

OpenStudy (koikkara):

yeah it does, you already have the answer as u said D= M/V D= 425g/48.0cm3 = 8.85g/cm3 It is just a hint to solve your question and approximately your answer comes close to copper just a point difference right ?

OpenStudy (koikkara):

@AndrewKaiser333

OpenStudy (koikkara):

Seems Like you practice Less your lessons, Well Download this booklet and practise till you solve within few minutes. Hey @AndrewKaiser333 You in here or somewhere ? Have A Nice Day ! Close this Question Too !

OpenStudy (andrewkaiser333):

sorry had to take a shower and was sick the day it was taught or i would know it already the migraines and vomiting has been kickin my butt. i am a very smart kid but if i am not there it is a little hard at first till someone helps. i am sorry